The market demand and supply for renting apartments in the Midrand area is given as:

Demand: p = 12 000 - 5Qd

Supply: p = 800 + 3Qs


5. What is the landlord’s total revenue at the equilibrium price?

[1] R7 000 000.

[2] R12 000 000.

[3] R8 000 000.

[4] R1 400 000.


6. If the government implements an effective rent ceiling of R4 100 in the Midrand area, what are the search costs?

[1] R0.

[2] 2 640 000.

[3] 1 320 000.

[4] 72 000.


7. What price will the customers be willing to pay on the black market?

[1] R5 000.

[2] R4 100.

[3] R6 500.

[4] R1 400.


1
Expert's answer
2021-05-27T10:58:21-0400

5)"Q_D=Q_S"


"12,000-5Q=800+3Q"

"12,000-800=5Q+3Q"

"11,200=8Q"

"Q=1,400"


"P=12,000-5Q"

"P=12,000-5(1400)"

"P=12,000-7000"

"P=5,000"


Total Revenue ="P\u00d7Q"

"=R5, 000\u00d71,400"

"=R. 7,000,000"


6)Willingness to pay =12,000

Price Charge =5,000

Difference "=12,000-5000"

"=7,000"


Search Costs ="7,000\u00d71,400"

"=9,800,000"


7)The price in the black market would be closer to the consumer's willingness to pay which would extract the entire consumer surplus thus price in the black market would be R6,500. The black market price would be above the legal ceiling resulting in price above 4,100.The option 3 is correct.